Ivanhoe Corporation earned net income of $275,800 in 2020 and had 116,000 shares of common stock outstanding throughout the year. Also outstanding all year was $900,000 of 9% bonds, which are convertible into 17,000 shares of common. Ivanhoe’s tax rate is 30 percent.

Compute Ivanhoe’s 2020 diluted earnings per share. (Round answer to 2 decimal places, e.g. 3.55.)

Diluted earnings per share $Entry field with incorrect answer
Answer #1

Basic earning per share = 275800/116000 = 2.38

Diluted earning per share = Adjusted net income/Adjusted shares

Adjusted net income = 275800+(900000*9%*70%) = 332500

Adjusted shares = 116000+17000 = 133000

Diluted earning per share = 2.38 per share

Diluted earning per share = 332500/133000 =2.50